### v1

# Movie2Story: A framework for understanding videos and telling stories in the form of novel text

Abstract
Multimodal video-to-text models have made considerable progress, primarily in generating brief descriptions of video content. However, there is still a deficiency in generating rich long-form text descriptions that integrate both video and audio. In this paper, we introduce a framework called M2S, designed to generate novel-length text by combining audio, video, and character recognition. M2S includes modules for video long-form text description and comprehension, audio-based analysis of emotion, speech rate, and character alignment, and visual-based character recognition alignment. By integrating multimodal information using the large language model GPT4o, M2S stands out in the field of multimodal text generation. We demonstrate the effectiveness and accuracy of M2S through comparative experiments and human evaluation. Additionally, the model framework has good scalability and significant potential for future research.
Movie2Story: A framework for understanding videos and telling stories in the form of novel text
Kangning Li kangningl25@ucla.edu                        Zheyang Jia u3621320@connect.hku.hk                        Anyu Ying anyuyay@gmail.com
1 Introduction
Many popular movies are inaccessible to disabled individuals who can only listen to subtitles, missing critical details such as environment and expressions, leading to a poor viewing experience. Additionally, people often prefer reading books to understand story development, as text is not constrained by movie length and provides greater convenience. Large language models (LLMs) have shown significant success in text understanding and generation, making them suitable for story comprehension and novel writing. Pre-trained multimodal models can map audio, images, and text spaces, enabling text generation when combined with LLMs. This integration can produce novels with detailed environmental and psychological descriptions, and a coherent storyline.We aim to develop a comprehensive video-language model that integrates multimodal information from video and audio to translate films into novels with complete elements. This model will cater to diverse groups, including barrier-free users and those who prefer e-books.
In the field of video generated text, multimodal understanding models based on CLIP/BLIP have made good progress (reference A). Some models focus on question and answer understanding for videos, also known as VQA (reference B.), but only provide brief descriptions. However, longer descriptions such as paper C still lack detailed information. Some models choose to use audio to add detailed information, such as paper D, but do not generate sufficiently long and rich text. In contrast, paper E chooses to generate long texts using a hierarchical structure, but lacks the ability to handle other multimodal processing such as audio.
Existing models can combine video and audio signals to generate text descriptions but lack deeper semantic information such as emotions and detailed environmental context. These models are primarily used for caption generation and question-answer scenarios, limiting their applicability to short film segments and not suitable for creating novels with rich elements.
In this article, our method combines the advantages of these papers by combining audio, video, and images to effectively generate novel level long texts with multimodal understanding capabilities. And these texts can be used as datasets for future researchers. Current video-language models focus primarily on generating captions, image descriptions, and question-answer scenes. While Our approach aims to leverage multimodal data, including audio, to extract richer information such as emotions, thereby creating detailed and engaging narratives from movies, referred to as movie2story. This method addresses the needs of disabled individuals and those who prefer reading, enhancing their ability to enjoy and understand movie content through comprehensive and immersive text.
By overcoming these challenges, our video-language model will generate detailed and immersive narratives from movies, enhancing accessibility and enriching the experience for all users.
This research developed new methodologies for combining audio, video, and textual data to generate comprehensive narratives. By advancing multimodal integration, we will provide significant insights into harmonizing different data types to produce richer, more detailed text. The enhanced model will be particularly relevant to disabled individuals who rely on text or audio descriptions to understand films, improving their movie-watching experience. Additionally, e-book readers who prefer reading will benefit from the comprehensive and immersive stories generated by the model.
Overall, our experimental results strongly demonstrate the ability of the model to generate text with complete elements and richer semantic information, making it adaptable to various application scenarios. It can be used in educational settings for detailed explanations of visual content or in entertainment to create novelizations of films for broader audiences.
2 Related Works
2.1 MMAD: Multi-modal Movie Audio DescriptionYe et al. (2024)
The MMAD (Multimodal Audio-Visual-Text Description) model combines video, audio, and textual information to achieve concise, real-time descriptions. This model leverages several key components to enhance the richness and accuracy of generated narratives:
By integrating these components, the MMAD model provides a comprehensive approach to generating rich, real-time descriptions that enhance accessibility and understanding for all users.
2.2 Distilling Vision-Language Models on Millions of VideosZhao et al. (2024)
This workflow introduces an effective method for transferring image-language models to video-language models to generate large amounts of high-quality video pseudo-labels. The detailed steps are as follows:
The generated pseudo-labels provide multi-granularity features, offering more semantic information and background description for our method:
2.3 VideoChat : Chat-Centric Video UnderstandingLi et al. (2024)
VideoChat-Text textualizes videos in stream. VideoChat-Embed encodes videos as embeddings. Both video content can be input in LLMs for multimodal understanding.
2.4 Video-LLaMA An Instruction-tuned Audio-Visual Language Model for Video UnderstandingZhang et al. (2023)
The ImageBind module, which has been pre-trained, aligns audio, video, and other multimodal information into a single vector space. Instead of training an audio-text dataset, it solely trains a video-text encoder, which indirectly allows the conversion of audio into text.
2.5 Video Storytelling: Textual Summaries for EventsLi et al. (2020)
In this paper, the proposed method consists of three main steps: key frame selection using the Narrator model, multi-modal embedding learning for contextual understanding, and story generation.
Key Frame Selection using Narrator Model: The Narrator model leverages reinforcement learning to select key frames from long videos. It determines the importance of each frame based on semantic similarity with previous frames and decides on the clip length:
Context-Aware Multi-Modal Embedding Learning: After selecting key frames, they use a bi-directional RNN (ResBRNN) to learn contextual embeddings that capture the semantic information of video frames and texts:
They propose replacing ResBRNN with Transformers to leverage their superior long-range dependency modeling capabilities.
This approach can significantly improve the performance and quality of video storytelling, making the generated stories more cohesive and diverse.
3 Preliminary
3.1 Datasets
InternVidLi et al. (2020) A high-quality, large-scale video-text dataset released in 2024 by the Shanghai AI Laboratory, Nanjing University, and the Chinese Academy of Sciences. InternVid contains over 7 million videos with detailed textual descriptions, covering 16 scenes and approximately 6000 action descriptions, with a total duration of nearly 760,000 hours. The videos and textual descriptions in this dataset are highly matched, making it suitable for video-text semantic matching, video-text retrieval, video-text generation, and other multimodal learning tasks.
VriptLi et al. (2020) Proposed by Shanghai Jiao Tong University, Beihang University, and the Xiaohongshu research team, the Vript dataset includes a meticulously annotated corpus of 12,000 high-resolution videos, providing detailed, dense, script-like subtitles for over 420,000 clips. Each clip’s subtitle contains approximately 145 words, which is ten times longer than most video-text datasets. Vript records not only content but also camera operations, including shot types and camera m
MSR-VTT (Microsoft Research Video to Text)Li et al. (2020) MSR-VTT is a large-scale video description dataset consisting of 10000 video clips divided into 20 categories. Each video clip is annotated with 20 English sentences by Amazon Mechanical Turks. Under standard conditions, the dataset is divided into a training set (6513 segments), a validation set (497 segments), and a testing set (2990 segments).
ActivityNet CaptionsLi et al. (2020) The ActiveNet Captions dataset is built on ActiveNet v1.3 and contains 20000 unedited YouTube videos with an average duration of 120 seconds and a total of 100000 subtitle annotations. Most videos contain more than three annotated events, each with a corresponding timestamp and manually written sentences. The dataset is divided into a training set (10024 videos), a validation set (4926 videos), and a testing set (5044 videos).
Video StoryLi et al. (2020) The Video Story dataset was created by Facebook to train AI systems to convert videos into stories. This dataset contains 20000 videos and 123000 descriptive sentences. These videos were selected from social media and have a high level of engagement, covering a wide range of topics and categories. The length of each video ranges from 20 to 180 seconds, with annotated paragraphs describing the object, context, and important details. The fragment contains an average of five sentences, with each sentence taking approximately 18 seconds on average.
3.2 Tools and Models
M2S maintains a Foundation Models Pool to store various video foundation models, which own the capability to detect and annotate different attributes of tracklets in videos. Here we list several models in it:
Video Frame Extraction Tool: FFmpeg, used for extracting key frames from videos.
Audio Extraction Tool: Whisper model for speech-to-text conversion, noting it generates text without capturing emotions.
Pre-trained Models: Including MMAD (video-language model) and VideoChat for generating detailed descriptions and dialogue content.
Emotion Analysis Module: Using features like speech intonation for audio-emotion processing, or BERT-like encoder networks for text emotion analysis to extract emotional information.
Character Relationship Network: To enable the model to maintain story coherence based on character relationships.
Contextual Information Linking: Utilizing resampling or contextual information to ensure semantic consistency.
Long Sequence Processing: Using batch processing and key frame resampling techniques to link information across different segments.
Large Language Models: Such as GPT-4, used to expand and refine generated descriptions into a complete novel.
4 Methodology
Figure 1: Workflow template
4.1 Image Character Recognition
Our objective is to achieve the matching of real-person names with characters appearing in the video. This necessitates the initial recognition of individuals in the footage, as well as the matching of the same individuals across consecutive video frames. Given that our application scenario involves long videos, we must contend with challenges such as multiple viewpoints of individuals and variations in their appearance, for instance, wearing different attire at various time intervals. These factors present significant challenges in the matching of individuals in extended video footage.
Drawing inspiration from the approach proposed in the MMAD paper, which transforms the problem into a person Re-identification (ReID) task, we utilized text feature information to enhance the spatial features of CLIP, thereby further distinguishing individual characteristics and achieving effective person identification and matching.
Considering that the appearance and behavior of individuals should remain consistent within shorter video segments, we can also employ person tracking algorithms for matching within these segments. However, given our ultimate application scenario, we still need to match the same individuals across different segments, which inevitably involves dealing with changes in appearance.
To address the person matching challenge, we have chosen to utilize Facenet, which can accurately identify faces in individual images and compute similarity for person matching. Since facial features do not change with the scene, this approach offers robustness. Additionally, we propose a text feature-based matching method for short video segments. This involves using VIT to generate descriptions of individuals’ appearances, which are then processed by LLM to determine the matching between different descriptions. By combining vector features with textual descriptions, we are able to effectively match individuals across various scenes in long videos.
For the global person-name matching, we employ an anchor matching method. We pre-establish a dictionary of key characters in the video, pairing names with ‘face features, textual appearance features’. Using these set anchors, we can allocate names to the detected individuals in the video.
Considering the angle of pre-obtained person bounding boxes, we propose using the YOLO method to pre-segment the bounding box information of individuals. Unprocessed images may contain multiple individuals, which can interfere with the extraction of facial and textual features. Therefore, we utilize YOLOv8 for the preprocessing and extraction of person bounding boxes. However, YOLO extracts all individuals from the frame image, so we subsequently determine whether a character is a main character by comparing the similarity between anchors and candidates, thereby filtering the information.
4.2 Video
Our approach employs a segmented video processing strategy, which effectively addresses the challenges of handling long videos while maintaining the continuity of the narrative. Within each video segment, we focus on extracting storyline information and background context as the key components of the video segment. In terms of model selection, we conducted a comprehensive evaluation of various open-source models, including videollava and Internvideo. Following a comparative analysis, we finally adopted for the open-source model Videochat as our baseline model. By carefully designing the prompt format, we ensured that the model could accurately extract the necessary key information, laying the foundation for in-depth analysis of the video content.
4.3 Audio
Our audio processing pipeline begins with the extraction of audio from video files using Python’s audio libraries. The extracted audio is then segmented into smaller, manageable chunks, facilitating efficient processing and analysis. The data is stored in JSON format for downstream text processing, which includes not only audio features and content but also timestamps. Therefore, it can be processed by subsequent LLMs to obtain highly consistent text content.
4.3.1 Speaker Recognition
Speaker recognition is an important branch of speech technology aimed at identifying or verifying the identity of a speaker based on their speech features. It extracts the unique identity information of the speaker by analyzing the acoustic features of the speech signal, and performs matching and recognition. Speaker recognition technology plays an important role in fields such as security monitoring, intelligent voice assistants, and identity authentication. In this article, we use distinguishing different speakers to achieve audio and video character matching for accurate text description.
1. Feature Extraction:
To capture the acoustic characteristics of each audio segment, we employ the pyannote.audio toolkit, which offers a range of feature extraction techniques. This includes standard methods like Mel-frequency cepstral coefficients (MFCCs) and spectrograms, providing a rich representation of the audio signal. Additionally, pyannote.audio supports on-the-fly data augmentation, enhancing the robustness and generalization of the models.
2. Speaker Embedding:
The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audio. This module leverages deep learning models to extract unique and discriminative features from each speaker’s voice. These speaker embeddings serve as compact representations that capture the essential aspects of a speaker’s identity, facilitating accurate recognition and matching.
pyannote.audio provides a unified framework for training sequence labeling models for tasks like Voice Activity Detection, Speaker Change Detection, and Overlapping Speech Detection. By employing recurrent neural networks and attention mechanisms, these models effectively learn to identify and segment different speakers within the audio stream.
To further improve speaker recognition performance, pyannote.audio employs metric learning techniques to train speaker embeddings that are optimized for a predefined distance metric, such as cosine similarity. This approach eliminates the need for additional steps like Probabilistic Linear Discriminant Analysis (PLDA), simplifying the clustering process.
3. End-to-End Training and Pipeline Integration:
pyannote.audio supports end-to-end training, allowing us to directly learn from the raw audio waveform without the need for manual feature extraction. This streamlines the workflow and enables the models to capture more nuanced acoustic information.
The pyannote.audio pipeline module integrates these modules into a cohesive framework, optimizing hyperparameters and jointly training the entire speaker segmentation pipeline. This approach, as opposed to independently training and combining individual modules, often leads to improved performance and more accurate speaker segmentation results.
4.3.2 Audio Description
In the Audio Description module, we leverage the powerful Whisper model to extract specific speech content from the audio segments. Whisper, an open-source speech recognition system developed by OpenAI, demonstrates exceptional performance across a wide range of languages and dialects. It excels in accurately transcribing speech, even in challenging acoustic environments. By employing end-to-end unsupervised training, Whisper efficiently learns from raw audio data, eliminating the need for manual annotations and significantly reducing training costs. Whisper’s versatility extends beyond speech recognition, as it can also be utilized for speech translation into multiple languages, enabling cross-lingual communication. Additionally, its ability to search for specific audio content based on keywords or phrases enhances its utility in information retrieval tasks. The modular design of Whisper, along with its availability in various sizes, allows for flexible deployment options based on specific needs and resource constraints. The open-source nature of Whisper further promotes collaboration and innovation within the speech recognition community, fostering the development of new applications and advancements in the field.
4.3.3 Emotion Analysis
Emotion2Vec is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space. This method not only captures the subtle relationships between words, but also assigns them the emotional burden of human expression.
The journey of Emotion2Vec begins with careful planning of the training dataset, where each word is carefully paired with its emotional essence - whether it’s discrete categories of happiness or anger, or subtle levels of intensity and arousal. From the perspective of models such as Word2Vec or GloVe, these words take on a new form, transformed into vectors that preserve their semantic structure while encoding their emotional meanings. As each emotional word gains its unique vector representation, a series of emotional vectors unfold, revealing the closeness of "happiness" to "joy" and its distance from "sadness". This mapping can provide a profound understanding of how emotions are integrated into language structures.
In the field of sentiment analysis, Emotion2Vec shines brightly, able to interpret emotions in various forms of data - whether it’s written text, spoken language, or visual images. It accurately classifies emotions, estimates their intensity, and identifies the most subtle similarities in emotions. The advantage of Emotion2Vec lies in its ability to grasp subtle semantic differences, express emotions clearly, and easily adapt to new emotional landscapes. Its applications cover the analysis of social media conversations, the classification of emotions in literary works, and the comparison of emotions between different datasets.
Ultimately, Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language.The model in this article is derived from the Emotion2Vec related model trained by FunASR.
4.3.4 Word Speed
3 This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow"
5 Experiment
5.1 Quantitative Analysis
Traditional NLP metrics such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015) evaluate generated text by comparing it to reference ground truth descriptions. While effective for assessing surface-level language quality, these metrics fail to capture the richness of knowledge beyond the visual domain, such as audio-related information, character details, or contextual nuances.
For instance, while a visual scene may depict a list of ingredients, a narration could highlight only a subset of those ingredients or incorporate additional elements like speaker emotions or environmental sounds, which are critical for generating a more comprehensive story. Since traditional metrics rely solely on visual references, they do not account for these extra layers of information.
To address this gap, we propose a set of reference-free metrics: Language Fluency, Key-knowledge Relevance. These metrics evaluate fluency, knowledge integration (including both visual and audio elements), and narrative alignment, providing a more accurate assessment of the generated stories and better reflecting the advantages of our multi-modal framework.
Language Fluency: To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story. Repetition of triplets within the story is considered a sign of unnatural phrasing and redundancy, which negatively impacts fluency. A higher fluency score is assigned when the text flows smoothly, with minimal repetition and a coherent structure.
Key-knowledge Relevance: For evaluating the relevance of key knowledge in the generated story, we define a knowledge base that includes both visual and audio information. In contrast to traditional models that focus solely on visual knowledge, our approach incorporates additional audio-related knowledge, such as ASR (Automatic Speech Recognition), environmental sound events, and other audio features. This expanded knowledge base allows us to assess the relevance and diversity of knowledge within the story. Specifically, we use two evaluation criteria: - Information Similarity (InfoSim): Measures the alignment between the knowledge points in the story and the knowledge repository. A higher similarity score indicates that the generated story effectively incorporates relevant knowledge. - Information Diversity (InfoDiverse): Evaluates the breadth of the knowledge used in the story. A higher diversity score indicates that the story incorporates a wide range of knowledge points, avoiding over-reliance on a small subset of information.
A well-rounded story should demonstrate both high information similarity and diversity, ensuring that relevant visual and audio knowledge are integrated in a meaningful way.
These evaluation metrics provide a comprehensive framework for assessing the quality of generated story descriptions, ensuring that the generated narratives meet high standards in terms of fluency, knowledge relevance, and visual coherence.
5.2 Qualitative Analysis
5.2.1 GPT Assistant Metrics
In this section, we introduce a set of key evaluation metrics designed to assess the quality of generated story descriptions using GPT and prompt engineering. By defining these metrics, we aim to leverage GPT as a rigorous evaluator of novel-like text, assigning scores based on specific story elements. These metrics are particularly useful for assessing narrative quality in terms of various aspects such as environmental details, emotional expression, language proficiency, character portrayal, and storyline coherence.
The following metrics are defined:
Environment Description Score: This score evaluates how well the generated text sets the scene and conveys the environment. Key factors include the richness of the setting, sensory detail, and atmosphere. A high score indicates a vivid and immersive portrayal of the environment.
Emotional Description Score: This metric assesses the depth and authenticity of emotional expression in the story. It measures how well the text captures and conveys the emotional states of characters, as well as the emotional tone of the narrative.
Language Description Score: This score focuses on the clarity, fluency, and grammatical quality of the text. It evaluates how well the story is written, including sentence structure, vocabulary choice, and coherence. A higher score reflects a more polished and readable narrative.
Character Description Score: This metric evaluates how effectively characters are described in terms of appearance, personality, and behavior. It gauges whether the character descriptions are detailed, nuanced, and consistent throughout the narrative.
Structural and Rationality Coherence Index (SRCI) Score: This metric quantifies the narrative coherence and the logical progression of the storyline. It gauges the effectiveness of plot construction, the lucidity of the story’s trajectory, and the seamless incorporation of essential storytelling components such as conflict, resolution, and character progression. The SRCI reflects the story’s ability to maintain a consistent and engaging structure that resonates with the audience’s expectations of narrative rationality.”
Storyline Accuracy with Canonical Outline Reference (SACOR) Score: This score measures the accuracy of the generated story against a predefined canonical storyline. It assesses how closely the generated narrative aligns with the key plot points, character behaviors, and thematic elements of the original story outline. The SACOR score is indicative of the model’s fidelity to the source material and its ability to reproduce the intended narrative arc with precision.”
For a more comprehensive evaluation, these scores are tabulated and compared across different story generations. The results can be found in Table 2, which summarizes the performance of our framework on each metric.
In Table 2, the scores represent the average evaluation across multiple generated stories, providing a comprehensive view of the performance of our framework compared to baseline methods.
5.2.2 Manual Metrics
Table 2 shows the results of applying our M2S framework for generating detailed story descriptions from several movies. The comparison between our framework and baseline methods demonstrates the clear advantages of our approach.
More specifically, our framework excels in producing more comprehensive and accurate story descriptions. By leveraging the Audio module, we incorporate key information such as ASR (Automatic Speech Recognition) output, speaker identity, and speaking rate to further enhance the generated descriptions. Additionally, the combination of the Actor-Matching module, Storyline-Segment module, Audio module, and Vision module enables our framework to generate rich, long-form narrative descriptions that go beyond traditional subtitles, offering a novel way of understanding movies. This integrated approach results in more detailed and accurate descriptions, which could serve as an independent means of movie comprehension, rather than merely being a supplementary subtitle track.
6 Conclusion
Acknowledgments
Appendix A Example Appendix
This is an appendix.
Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
Internvideo2 61.98 38.00 11.24 0.257 0.295 0.108 0.0265 0.0154
VideoLlava 15.94 38.18 13.64 0.376 0.275 0.113 0.0334 0.00474
M2s(GPT4o) 10.24 30.10 59.11 0.175 0.148 0.0556 0.0128 1.25e-6
M2s(llama) - - -
M2s(Flan-T5) - - -
Table 1: Performance of various models on Quantitative metrics
Model Environment Language Emotion Character SRCI SACOR_T Overall GPT Human
Internvideo2 1.0 1.0 0.0 0.6 0.2 0.600 0.56 -
VideoLlava 1.1 1.2 0.0 0.9 0.9 0.559 0.8 -
M2s(GPT4o) 2.7 2.3 1.0 1.7 1.7 0.735 1.9 -
M2s(llama) - - - - - -
M2s(Flan-T5) - - - - - -
Table 2: Performance of various models on Qualitative metrics
Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
M2s(without emo&speed) 7.31 29.87 52.26 0.196 0.166 0.0651 0.0137 1.24e-6
M2s(without ASR&emo&speed) 14.58 29.43 37.12 0.199 0.140 0.0626 0.0217 0.00365
M2s(GPT4o) 10.24 28.63 83.30 0.175 0.148 0.0556 0.0128 1.25e-6
Table 3: Performance of Ablation Study

### v2

@@
 Abstract
-Multimodal video-to-text models have made considerable progress, primarily in generating brief descriptions of video content.
-However, there is still a deficiency in generating rich long-form text descriptions that integrate both video and audio.
-In this paper, we introduce a framework called M2S, designed to generate novel-length text by combining audio, video, and character recognition.
-M2S includes modules for video long-form text description and comprehension, audio-based analysis of emotion, speech rate, and character alignment, and visual-based character recognition alignment.
-By integrating multimodal information using the large language model GPT4o, M2S stands out in the field of multimodal text generation.
-We demonstrate the effectiveness and accuracy of M2S through comparative experiments and human evaluation.
-Additionally, the model framework has good scalability and significant potential for future research.
+In recent years, large-scale models have achieved significant advancements, accompanied by the emergence of numerous high-quality benchmarks for evaluating various aspects of their comprehension abilities.
+However, most existing benchmarks primarily focus on spatial understanding in static image tasks.
+While some benchmarks extend evaluations to temporal tasks, they fall short in assessing text generation under complex contexts involving long videos and rich auxiliary information.
+To address this limitation, we propose a novel benchmark: the Multi-modal Story Generation Benchmark (MSBench), designed to evaluate text generation capabilities in scenarios enriched with auxiliary information.
+Our work introduces an innovative automatic dataset generation method to ensure the availability of accurate auxiliary information.
+On one hand, we leverage existing datasets and apply automated processes to generate new evaluation datasets, significantly reducing manual efforts.
+On the other hand, we refine auxiliary data through systematic filtering and utilize state-of-the-art models to ensure the fairness and accuracy of the ground-truth datasets.
+Our experiments reveal that current Multi-modal Large Language Models (MLLMs) perform suboptimally under the proposed evaluation metrics, highlighting significant gaps in their capabilities.
+To address these challenges, we propose a novel model architecture and methodology to better handle the overall process, demonstrating improvements on our benchmark.
 Movie2Story: A framework for understanding videos and telling stories in the form of novel text
-Kangning Li kangningl25@ucla.edu                        Zheyang Jia u3621320@connect.hku.hk                        Anyu Ying anyuyay@gmail.com
+Kangning Li kangningl25@ucla.edu                        Zheyang Jia u3621320@connect.hku.hk                        Anyu Ying anyuy@andrew.cmu.edu
 1 Introduction
-Many popular movies are inaccessible to disabled individuals who can only listen to subtitles, missing critical details such as environment and expressions, leading to a poor viewing experience.
-Additionally, people often prefer reading books to understand story development, as text is not constrained by movie length and provides greater convenience.
-Large language models (LLMs) have shown significant success in text understanding and generation, making them suitable for story comprehension and novel writing.
-Pre-trained multimodal models can map audio, images, and text spaces, enabling text generation when combined with LLMs.
-This integration can produce novels with detailed environmental and psychological descriptions, and a coherent storyline.We aim to develop a comprehensive video-language model that integrates multimodal information from video and audio to translate films into novels with complete elements.
-This model will cater to diverse groups, including barrier-free users and those who prefer e-books.
-In the field of video generated text, multimodal understanding models based on CLIP/BLIP have made good progress (reference A).
-Some models focus on question and answer understanding for videos, also known as VQA (reference B.), but only provide brief descriptions.
-However, longer descriptions such as paper C still lack detailed information.
-Some models choose to use audio to add detailed information, such as paper D, but do not generate sufficiently long and rich text.
-In contrast, paper E chooses to generate long texts using a hierarchical structure, but lacks the ability to handle other multimodal processing such as audio.
-Existing models can combine video and audio signals to generate text descriptions but lack deeper semantic information such as emotions and detailed environmental context.
-These models are primarily used for caption generation and question-answer scenarios, limiting their applicability to short film segments and not suitable for creating novels with rich elements.
-In this article, our method combines the advantages of these papers by combining audio, video, and images to effectively generate novel level long texts with multimodal understanding capabilities.
-And these texts can be used as datasets for future researchers.
-Current video-language models focus primarily on generating captions, image descriptions, and question-answer scenes.
-While Our approach aims to leverage multimodal data, including audio, to extract richer information such as emotions, thereby creating detailed and engaging narratives from movies, referred to as movie2story.
-This method addresses the needs of disabled individuals and those who prefer reading, enhancing their ability to enjoy and understand movie content through comprehensive and immersive text.
-By overcoming these challenges, our video-language model will generate detailed and immersive narratives from movies, enhancing accessibility and enriching the experience for all users.
-This research developed new methodologies for combining audio, video, and textual data to generate comprehensive narratives.
-By advancing multimodal integration, we will provide significant insights into harmonizing different data types to produce richer, more detailed text.
-The enhanced model will be particularly relevant to disabled individuals who rely on text or audio descriptions to understand films, improving their movie-watching experience.
-Additionally, e-book readers who prefer reading will benefit from the comprehensive and immersive stories generated by the model.
-Overall, our experimental results strongly demonstrate the ability of the model to generate text with complete elements and richer semantic information, making it adaptable to various application scenarios.
-It can be used in educational settings for detailed explanations of visual content or in entertainment to create novelizations of films for broader audiences.
+In recent years, many multi-modal large language models (MLLMs)Alayrac et al.
+(2022); Zhu et al.
+(2023); Li et al.
+(2024a); Huang et al.
+(2023); Li et al.
+(2020) have effectively combined video encodersTouvron et al.
+(2023); Devlin et al.
+(2019); Dosovitskiy et al.
+(2021) and audio encodersRadford et al.
+(2022); Chen et al.
+(2022b, a) to achieve strong performance in vision-based description generation and audio-based tasks.
+However, as MLLMs evolve, a critical challenge arises: how can we rigorously evaluate their comprehension and generation capabilities?
+Addressing this question is vital for assessing current models and guiding future developments.
+Existing benchmarks for MLLMs primarily adopt a question-answering (QA) formatYu et al.
+(2024); Xu et al.
+(2023); Xie et al.
+(2024); Fu et al.
+(2024a); Xu et al.
+(2016); Liu et al.
+(2024); Cheng et al.
+(2024), focusing on static image understanding.
+While benchmarks such as MVBenchLi et al.
+(2024b) extend evaluation to temporal tasks, they remain inadequate for assessing long-video comprehension and fail to incorporate rich auxiliary information such as audio.
+Although some studies, such as AV-SUPERBTseng et al.
+(2024), attempt to integrate audio information with visual information, they primarily focus on the evaluation of audio.
+Moreover, current benchmarks typically produce objective, template-like outputs, lacking the stylistic complexity of narrative storytelling.
+Furthermore, many benchmarks rely heavily on manual annotations, which are resource-intensive and time-consuming.
+To overcome these limitations, we introduce a novel benchmark, Multi-modal Story Generation Benchmark (MSBench), which emphasizes the integration of diverse auxiliary information (e.g., audio features) to generate long-video descriptions in a narrative and information-rich style.
+Our approach features an innovative data transformation pipeline tailored to enhance existing video-text datasetsLi et al.
+(2020).
+Recognizing the brevity and limited scope of current datasets, we augment them with detailed audio and visual information.
+For instance, simple descriptions like "a car is driving" are enriched into nuanced narratives such as "a man sings: ‘…’, while a woman drives the car…".
+Auxiliary information, including audio-derived attributes such as ASR (Automatic Speech Recognition)Kheddar et al.
+(2024) outputs, emotions, and sound events, is extracted using powerful open-source audio models.
+For visual descriptions, we leverage state-of-the-art image understanding modelsBai et al.
+(2023); Chen et al.
+(2024) to generate detailed frame-level descriptions, ensuring narrative consistency through temporal ordering.
+Our benchmark offers two significant advantages.
+First, it drastically reduces the reliance on manual annotation by utilizing automated processes and open-source tools.
+Second, it introduces five novel evaluation metrics tailored to assess text quality in enriched information scenarios.
+These metrics span dimensions from environmental and psychological descriptions to name-entity specificity, providing a comprehensive assessment of narrative text quality in open-world, long-duration contexts.
+Finally, we evaluate several state-of-the-art MLLMsLi et al.
+(2024a); Lin et al.
+(2024); Wang et al.
+(2022, 2024) on the MSBench benchmark, revealing substantial gaps in their performance.
+Models like VideoChat2Li et al.
+(2024b) struggle with incorporating audio features, handling long-video contexts, and generating high-quality narrative outputs.
+To address these challenges, we propose a robust baseline model, M2S-LLM, specifically designed to align with our benchmark’s requirements and enhance text generation quality.
+Experimental results demonstrate that M2S-LLM surpasses VideoChat2Li et al.
+(2024b) by nearly 15% across key metrics.
+All models, datasets, and evaluation frameworks are publicly available to facilitate future research and advancements in the field.
 2 Related Works
-2.1 MMAD: Multi-modal Movie Audio DescriptionYe et al. (2024)
-The MMAD (Multimodal Audio-Visual-Text Description) model combines video, audio, and textual information to achieve concise, real-time descriptions.
-This model leverages several key components to enhance the richness and accuracy of generated narratives:
-By integrating these components, the MMAD model provides a comprehensive approach to generating rich, real-time descriptions that enhance accessibility and understanding for all users.
-2.2 Distilling Vision-Language Models on Millions of VideosZhao et al. (2024)
-This workflow introduces an effective method for transferring image-language models to video-language models to generate large amounts of high-quality video pseudo-labels.
-The detailed steps are as follows:
-The generated pseudo-labels provide multi-granularity features, offering more semantic information and background description for our method:
-2.3 VideoChat : Chat-Centric Video UnderstandingLi et al. (2024)
-VideoChat-Text textualizes videos in stream. VideoChat-Embed encodes videos as embeddings. Both video content can be input in LLMs for multimodal understanding.
-2.4 Video-LLaMA An Instruction-tuned Audio-Visual Language Model for Video UnderstandingZhang et al. (2023)
-The ImageBind module, which has been pre-trained, aligns audio, video, and other multimodal information into a single vector space.
-Instead of training an audio-text dataset, it solely trains a video-text encoder, which indirectly allows the conversion of audio into text.
-2.5 Video Storytelling: Textual Summaries for EventsLi et al. (2020)
-In this paper, the proposed method consists of three main steps: key frame selection using the Narrator model, multi-modal embedding learning for contextual understanding, and story generation.
-Key Frame Selection using Narrator Model: The Narrator model leverages reinforcement learning to select key frames from long videos.
-It determines the importance of each frame based on semantic similarity with previous frames and decides on the clip length:
-Context-Aware Multi-Modal Embedding Learning: After selecting key frames, they use a bi-directional RNN (ResBRNN) to learn contextual embeddings that capture the semantic information of video frames and texts:
-They propose replacing ResBRNN with Transformers to leverage their superior long-range dependency modeling capabilities.
-This approach can significantly improve the performance and quality of video storytelling, making the generated stories more cohesive and diverse.
-3 Preliminary
-3.1 Datasets
-InternVidLi et al.
-(2020) A high-quality, large-scale video-text dataset released in 2024 by the Shanghai AI Laboratory, Nanjing University, and the Chinese Academy of Sciences.
-InternVid contains over 7 million videos with detailed textual descriptions, covering 16 scenes and approximately 6000 action descriptions, with a total duration of nearly 760,000 hours.
-The videos and textual descriptions in this dataset are highly matched, making it suitable for video-text semantic matching, video-text retrieval, video-text generation, and other multimodal learning tasks.
-VriptLi et al.
-(2020) Proposed by Shanghai Jiao Tong University, Beihang University, and the Xiaohongshu research team, the Vript dataset includes a meticulously annotated corpus of 12,000 high-resolution videos, providing detailed, dense, script-like subtitles for over 420,000 clips.
-Each clip’s subtitle contains approximately 145 words, which is ten times longer than most video-text datasets.
-Vript records not only content but also camera operations, including shot types and camera m
-MSR-VTT (Microsoft Research Video to Text)Li et al.
-(2020) MSR-VTT is a large-scale video description dataset consisting of 10000 video clips divided into 20 categories.
-Each video clip is annotated with 20 English sentences by Amazon Mechanical Turks.
-Under standard conditions, the dataset is divided into a training set (6513 segments), a validation set (497 segments), and a testing set (2990 segments).
-ActivityNet CaptionsLi et al.
-(2020) The ActiveNet Captions dataset is built on ActiveNet v1.3 and contains 20000 unedited YouTube videos with an average duration of 120 seconds and a total of 100000 subtitle annotations.
-Most videos contain more than three annotated events, each with a corresponding timestamp and manually written sentences.
-The dataset is divided into a training set (10024 videos), a validation set (4926 videos), and a testing set (5044 videos).
-Video StoryLi et al.
-(2020) The Video Story dataset was created by Facebook to train AI systems to convert videos into stories.
-This dataset contains 20000 videos and 123000 descriptive sentences.
-These videos were selected from social media and have a high level of engagement, covering a wide range of topics and categories.
-The length of each video ranges from 20 to 180 seconds, with annotated paragraphs describing the object, context, and important details.
-The fragment contains an average of five sentences, with each sentence taking approximately 18 seconds on average.
-3.2 Tools and Models
-M2S maintains a Foundation Models Pool to store various video foundation models, which own the capability to detect and annotate different attributes of tracklets in videos.
-Here we list several models in it:
-Video Frame Extraction Tool: FFmpeg, used for extracting key frames from videos.
-Audio Extraction Tool: Whisper model for speech-to-text conversion, noting it generates text without capturing emotions.
-Pre-trained Models: Including MMAD (video-language model) and VideoChat for generating detailed descriptions and dialogue content.
-Emotion Analysis Module: Using features like speech intonation for audio-emotion processing, or BERT-like encoder networks for text emotion analysis to extract emotional information.
-Character Relationship Network: To enable the model to maintain story coherence based on character relationships.
-Contextual Information Linking: Utilizing resampling or contextual information to ensure semantic consistency.
-Long Sequence Processing: Using batch processing and key frame resampling techniques to link information across different segments.
-Large Language Models: Such as GPT-4, used to expand and refine generated descriptions into a complete novel.
-4 Methodology
+2.1 MLLM Introduction
+The evolution of Large Language Models (LLMs) has accelerated research into Multi-modal Large Language Models (MLLMs)Lyu et al.
+(2023); Lee et al.
+(2024); Fu et al.
+(2024b), aiming to integrate diverse modalities like text, vision, and audio.
+Early breakthroughs, such as FlamingoAlayrac et al.
+(2022) and PaLM-EDriess et al.
+(2023), showcased strong performance in multi-modal tasks by combining visual and textual modalities.
+Subsequent open-sourced efforts, including LLaVALiu et al.
+(2023) and MiniGPT-4Zhu et al.
+(2023), have expanded the scope of multi-modal instruction tuning, while VideoChatLi et al.
+(2024a) and VideoChatGPTMaaz et al.
+(2024) extended these ideas to dynamic video tasks by utilizing ChatGPT-generated annotations.
+Building on these foundations, several recent works have introduced innovative approaches to tackle the complexities of video comprehension.
+The MMADYe et al.
+(2024b) model integrates video, audio, and textual information to generate concise, real-time descriptions.
+By leveraging pre-trained encoders and advanced fusion techniques, MMADYe et al.
+(2024b) enhances narrative richness and accessibility.
+Similarly, the framework for Distilling Vision-Language Models on Millions of VideosZhao et al.
+(2024) adapts image-language models for video-language tasks, generating large-scale pseudo-labels with multi-granularity features, offering enhanced semantic and contextual understanding.
+The Video-LLaMA2Cheng et al.
+(2024) model aligns audio, video, and textual data into a unified vector space using the pre-trained ImageBindGirdhar et al.
+(2023) module.
+Instead of training an audio-text dataset, it utilizes a video-text encoder, indirectly converting audio into text.
+Meanwhile, Video StorytellingLi et al.
+(2020) focuses on generating textual summaries for events by selecting key frames with a reinforcement learning-based Narrator model and employing contextual embeddings through Residual Bidirectional RNN (ResBRNN)Li et al.
+(2020), with Transformers enhancing long-range dependency modeling for richer, more cohesive stories.
+Our work builds upon these advancements by addressing the limitations in existing MLLMs, particularly in handling long-duration videos, integrating auxiliary information, and generating stylistically rich narrative outputs.
+By combining innovative dataset enhancement techniques and robust evaluation metrics, our framework sets a new standard for multi-modal understanding and narrative-driven generation.
+2.2 Benchmark Introduction
+Traditional Vision-Language (VL) benchmarksGoyal et al.
+(2017); Kay et al.
+(2017); Xu et al.
+(2016, 2017); Xiao et al.
+(2021) have predominantly focused on QA-style evaluations, targeting tasks like multi-modal retrieval and vision QA.
+More recent benchmarks have broadened the scope to assess integrated capabilities.
+For example, OwlEvalYe et al.
+(2024a) and SEED-BenchLi et al.
+(2023) introduced evaluation metrics that transcend simple model ranking, focusing on comprehensive multi-modal skills.
+In the video domain, benchmarks like Perception TestPătrăucean et al.
+(2023) examine multi-modal video perception and reasoning, while FunQAXie et al.
+(2024) pushes reasoning boundaries using humorous or counter-intuitive content.
+MVBenchLi et al.
+(2024b) stands out by defining over 20 tasks to evaluate MLLMs’ performance on diverse scenarios, especially temporal reasoning.
+The study on Synchronized Video StorytellingYang et al.
+(2024) demonstrates an innovative approach by utilizing additional ad keywords for generating and assessing the quality of storytelling and advertisement content, providing considerable inspiration for the field.
+Similarly, MMADYe et al.
+(2024b) highlights the evaluation of critical but non-automatically measurable elements, such as character names, enriching benchmark design.
+These advancements inform our framework, which aims to address gaps in evaluating narrative-driven, multi-modal tasks.
+3 MSBench
+In this section, we delve into the details of our MSBench.
+Initially, we design the multi-modal story generation tasks as illustrated in Figure 1.
+Subsequently, we automatically generate caption-story pairs for evaluation, as depicted in stage 2 and outputs in Figure 1.
+3.1 Story Task Definition
+To craft the story tasks for MSBench, we employ a text-to-text methodology, which involves integrating scattered text information into long texts in the form of novels.
+As previously highlighted in the introduction, the majority of current MLLM benchmarks are centered around direct transformations from video to text or audio to text, often resulting in brief summary captions.
+Driven by this observation, we propose leveraging these existing task definitions as a foundation to systematically develop multi-modal story generation tasks.
+We begin by summarizing the main tasks of video understanding and audio understanding text summaries from previous benchmarks.
+Then we enrich these tasks to longer stories, creating story tasks that can not be effectively solved with a single model (video or audio) and require comprehensive understanding of video and audio information.
+Finally, we define 5 tasks as follows.
+video captions. A comprehensive understanding for video information, generated by VQA model like videochat2Li et al. (2024b).
+Audio ASR. Text version of the character’s speech in the audio, exactly what they said in the audio.
+Audio Feature. Emotion and speech speed summary for characters. And speaker recognition.
+Video Audio Alignment. Matching characters in movie.
+Timestamp matching and Story line Processing. Timestamp alignment for video and audio information and generation of storyline.
+3.2 Automatic caption-story pairs Generation
+M2S maintains a Foundation Models Pool to store various video and audio foundation models, which own the capability to detect and annotate different features in videos.
+After having sufficient features, we use LLM to integrate the features in chronological order and generate a lengthy novel level story text.
+M2S maintains a robust Foundation Models Pool.
+These models are equipped with the advanced capability to detect and annotate a myriad of features within videos.
+Once a sufficient array of features has been collected, we leverage a Large Language Model (LLM) to systematically integrate these features in chronological order (sorted by "timestamp" or segments as shown in Figure 1 video/audio captions respectively), thereby generating a comprehensive, novel-level story text that is both lengthy and rich in detail.
+3.2.1 Data Pre-processing
+(1) Video splitting: The video may have a long length, and a short 20s video is enough to generate a long story text.
+Therefore, dividing a long video into 20s has the following advantages: 1.
+Increasing the amount of data 2.
+Unifying the data size 3.
+Reducing memory usage.
+(2) FFmpeg: A Video Frame Extraction Tool, used for extracting key frames from videos.FFmpeg Developers (2023)
+3.2.2 Video-Audio Feature Information Generation..
+In our pipeline, feature extraction can be separate to several parts, here we list these models for extracting desired features and generating corresponding texts for our story tasks proposed in previous subsection.
+video captions. Use Pre-trained Models(video-language model), including MMADYe et al. (2024b) and VideoChat2Li et al. (2024b) for generating detailed descriptions and dialogue content.
+Audio ASR. Use speech Extraction Tool, eg. WhisperOpenAI (2022) model for speech-to-text conversion, noting it generates text without capturing emotions.
+Audio Feature. Use Emotion Analysis Module Using BERT-like encoder networks for audio emotion analysis to extract emotional information and transform in text version, eg Emotion2Vec.et al. (2023c)
+Video Audio Alignment.
+Character Relationship Network.
+Use CNNs like YOLOv8Yaseen (2024) for character recognition and audio processing models (eg.pyannote)et al.
+(2019) to split different characters.
+Matching them by anchors predefined by human.
+Timestamp matching and Story line Processing. Using audio timestamp splitting and video key frame resampling techniques to link information across different segments.
+3.2.3 Story Generation:
+Using Large Language Models, such as GPT-4oet al.
+(2024b), used to expand and refine generated descriptions into a complete novel.
+Here we use llama3-3Bet al.
+(2024a), mistral7Bet al.
+(2023a), Qwen7Bet al.
+(2023b) Collect all information and time sequence of video-audio text descriptions and generate a coherent story by well designed instructions and prompts.
+3.3 Prompt Design for Evaluation
+To emphasize the story line coherence of MLLMs, we craft a detailed system prompt for evaluation (see the top right of Figure 1).
+This prompt encourages MLLMs to carefully scrutinize video and audio content to generate stories, by paying attention to factors such as timestamps.
+3.4 Additional caption-stroy pair for fine-tuning LLMs
+We use GPTZhu et al.
+(2023) to generate virtual data in the corresponding format to enhance data quality.
+For example, as shown in Figure 1, first provide the format and example of the corresponding data for GPT.
+Then let GPT generate data in the corresponding format as virtual data and add it to the dataset for LLMs to make fine adjustments.
+Because at this stage, we only need to adjust the long text generated in the form of a novel, focusing mainly on whether the generated story contains the information in the caption and temporal alignment, rather than whether the information in the caption is in the original video and audio.
+Therefore, the data generated in this way is reasonable.
+4 M2S pipeline
 Figure 1: Workflow template
-4.1 Image Character Recognition
-Our objective is to achieve the matching of real-person names with characters appearing in the video.
-This necessitates the initial recognition of individuals in the footage, as well as the matching of the same individuals across consecutive video frames.
-Given that our application scenario involves long videos, we must contend with challenges such as multiple viewpoints of individuals and variations in their appearance, for instance, wearing different attire at various time intervals.
-These factors present significant challenges in the matching of individuals in extended video footage.
-Drawing inspiration from the approach proposed in the MMAD paper, which transforms the problem into a person Re-identification (ReID) task, we utilized text feature information to enhance the spatial features of CLIP, thereby further distinguishing individual characteristics and achieving effective person identification and matching.
-Considering that the appearance and behavior of individuals should remain consistent within shorter video segments, we can also employ person tracking algorithms for matching within these segments.
-However, given our ultimate application scenario, we still need to match the same individuals across different segments, which inevitably involves dealing with changes in appearance.
-To address the person matching challenge, we have chosen to utilize Facenet, which can accurately identify faces in individual images and compute similarity for person matching.
-Since facial features do not change with the scene, this approach offers robustness.
-Additionally, we propose a text feature-based matching method for short video segments.
-This involves using VIT to generate descriptions of individuals’ appearances, which are then processed by LLM to determine the matching between different descriptions.
-By combining vector features with textual descriptions, we are able to effectively match individuals across various scenes in long videos.
-For the global person-name matching, we employ an anchor matching method.
-We pre-establish a dictionary of key characters in the video, pairing names with ‘face features, textual appearance features’.
-Using these set anchors, we can allocate names to the detected individuals in the video.
-Considering the angle of pre-obtained person bounding boxes, we propose using the YOLO method to pre-segment the bounding box information of individuals.
-Unprocessed images may contain multiple individuals, which can interfere with the extraction of facial and textual features.
-Therefore, we utilize YOLOv8 for the preprocessing and extraction of person bounding boxes.
-However, YOLO extracts all individuals from the frame image, so we subsequently determine whether a character is a main character by comparing the similarity between anchors and candidates, thereby filtering the information.
-4.2 Video
-Our approach employs a segmented video processing strategy, which effectively addresses the challenges of handling long videos while maintaining the continuity of the narrative.
+After building our MSBench, we evaluate a number of popular video caption MLLMs in the first three rows in Table 2.
+Surprisingly, the existing MLLMs are far from satisfactory in temporal understanding.
+To fill the gap, we develop a robust MLLM-LLM pipeline, which is dubbed as M2S(movie to story).
+The basic process is the same as the way MSBench generates data.
+The basic process is the same as the way MSBench generates data.
+Here are some details of Video Audio Feature Information Generation and LLMs.
+The text integrated through LLM generally has a lower repetition rate and higher information score (column 2 and 3 in Table 2.)
+4.1 Video Feature Information Generation
+Our video processing pipeline first segments the video length, typically 20 seconds, to facilitate efficient processing and analysis.
+Then use open-source VLM and other models to extract video/image text descriptions from video files.
+The data is stored in JSON format for downstream text processing, which includes not only video content but also timestamps.
+Therefore, subsequent LLM can process it to obtain highly consistent text content.
+Video captions Our approach employs a segmented video processing strategy, which effectively addresses the challenges of handling long videos while maintaining the continuity of the narrative.
 Within each video segment, we focus on extracting storyline information and background context as the key components of the video segment.
-In terms of model selection, we conducted a comprehensive evaluation of various open-source models, including videollava and Internvideo.
-Following a comparative analysis, we finally adopted for the open-source model Videochat as our baseline model.
-By carefully designing the prompt format, we ensured that the model could accurately extract the necessary key information, laying the foundation for in-depth analysis of the video content.
-4.3 Audio
+In terms of model selection, we conducted a comprehensive evaluation of various open-source models, including videollavaLin et al.
+(2024) and InternvideoWang et al.
+(2022) .
+Following a comparative analysis, we finally adopted for the open-source model Videochat2Li et al.
+(2024b) as our baseline model.
+By carefully designing the prompt format for VQA models, we ensured that the model could accurately extract the necessary key information, laying the foundation for in-depth analysis of the video content.
+Image Character Recognition Our goal is to match real-person names with video characters by recognizing and matching individuals across frames in long videos, despite challenges like varying appearances and multiple viewpoints.
+We use FacenetSchroff et al.
+(2015) for robust facial recognition and a text feature-based method with VITDosovitskiy et al.
+(2021) and LLM for short-segment matching.
+For global matching, we establish a dictionary of key characters with face and text features as anchors.
+YOLOv8Yaseen (2024) is used for pre-segmenting person bounding boxes, and we filter main characters by comparing anchor-candidate similarity.
+This approach combines vector and text features to effectively match individuals across scenes in long videos.
+4.2 Audio Feature Information Generation
 Our audio processing pipeline begins with the extraction of audio from video files using Python’s audio libraries.
@@
 Therefore, it can be processed by subsequent LLMs to obtain highly consistent text content.
-4.3.1 Speaker Recognition
-Speaker recognition is an important branch of speech technology aimed at identifying or verifying the identity of a speaker based on their speech features.
-It extracts the unique identity information of the speaker by analyzing the acoustic features of the speech signal, and performs matching and recognition.
-Speaker recognition technology plays an important role in fields such as security monitoring, intelligent voice assistants, and identity authentication.
-In this article, we use distinguishing different speakers to achieve audio and video character matching for accurate text description.
-1. Feature Extraction:
-To capture the acoustic characteristics of each audio segment, we employ the pyannote.audio toolkit, which offers a range of feature extraction techniques.
-This includes standard methods like Mel-frequency cepstral coefficients (MFCCs) and spectrograms, providing a rich representation of the audio signal.
-Additionally, pyannote.audio supports on-the-fly data augmentation, enhancing the robustness and generalization of the models.
-2. Speaker Embedding:
-The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audio.
+Speaker Embedding for speaker recognition The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audioet al.
+(2019).
 This module leverages deep learning models to extract unique and discriminative features from each speaker’s voice.
 These speaker embeddings serve as compact representations that capture the essential aspects of a speaker’s identity, facilitating accurate recognition and matching.
-pyannote.audio provides a unified framework for training sequence labeling models for tasks like Voice Activity Detection, Speaker Change Detection, and Overlapping Speech Detection.
-By employing recurrent neural networks and attention mechanisms, these models effectively learn to identify and segment different speakers within the audio stream.
-To further improve speaker recognition performance, pyannote.audio employs metric learning techniques to train speaker embeddings that are optimized for a predefined distance metric, such as cosine similarity.
-This approach eliminates the need for additional steps like Probabilistic Linear Discriminant Analysis (PLDA), simplifying the clustering process.
-3. End-to-End Training and Pipeline Integration:
-pyannote.audio supports end-to-end training, allowing us to directly learn from the raw audio waveform without the need for manual feature extraction.
-This streamlines the workflow and enables the models to capture more nuanced acoustic information.
-The pyannote.audio pipeline module integrates these modules into a cohesive framework, optimizing hyperparameters and jointly training the entire speaker segmentation pipeline.
-This approach, as opposed to independently training and combining individual modules, often leads to improved performance and more accurate speaker segmentation results.
-4.3.2 Audio Description
-In the Audio Description module, we leverage the powerful Whisper model to extract specific speech content from the audio segments.
-Whisper, an open-source speech recognition system developed by OpenAI, demonstrates exceptional performance across a wide range of languages and dialects.
-It excels in accurately transcribing speech, even in challenging acoustic environments.
-By employing end-to-end unsupervised training, Whisper efficiently learns from raw audio data, eliminating the need for manual annotations and significantly reducing training costs.
-Whisper’s versatility extends beyond speech recognition, as it can also be utilized for speech translation into multiple languages, enabling cross-lingual communication.
-Additionally, its ability to search for specific audio content based on keywords or phrases enhances its utility in information retrieval tasks.
-The modular design of Whisper, along with its availability in various sizes, allows for flexible deployment options based on specific needs and resource constraints.
-The open-source nature of Whisper further promotes collaboration and innovation within the speech recognition community, fostering the development of new applications and advancements in the field.
-4.3.3 Emotion Analysis
-Emotion2Vec is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space.
+Emotion Analysis Emotion2Vecet al.
+(2023c) is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space.
 This method not only captures the subtle relationships between words, but also assigns them the emotional burden of human expression.
-The journey of Emotion2Vec begins with careful planning of the training dataset, where each word is carefully paired with its emotional essence - whether it’s discrete categories of happiness or anger, or subtle levels of intensity and arousal.
-From the perspective of models such as Word2Vec or GloVe, these words take on a new form, transformed into vectors that preserve their semantic structure while encoding their emotional meanings.
-As each emotional word gains its unique vector representation, a series of emotional vectors unfold, revealing the closeness of "happiness" to "joy" and its distance from "sadness".
-This mapping can provide a profound understanding of how emotions are integrated into language structures.
-In the field of sentiment analysis, Emotion2Vec shines brightly, able to interpret emotions in various forms of data - whether it’s written text, spoken language, or visual images.
-It accurately classifies emotions, estimates their intensity, and identifies the most subtle similarities in emotions.
-The advantage of Emotion2Vec lies in its ability to grasp subtle semantic differences, express emotions clearly, and easily adapt to new emotional landscapes.
-Its applications cover the analysis of social media conversations, the classification of emotions in literary works, and the comparison of emotions between different datasets.
-Ultimately, Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language.The model in this article is derived from the Emotion2Vec related model trained by FunASR.
-4.3.4 Word Speed
-3 This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow"
+Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language.
+The model in this article is derived from the Emotion2Vec related model trained by FunASR.
+Word Speed This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow".
+4.3 LLMs selection and LoRA Finetuning
+Different LLMs have varying abilities to integrate information and respond differently to instructions.
+Moreover, as a proprietary task, generic LLM may not necessarily have good results, so techniques such as LoRA may be needed to fine tune it.
+After extracting video and audio information, we can fine tune LoRAet al.
+(2021) for different LLMs to complete the final stage of generation and obtain stories.
+Note that the data here can be real video audio caption story pairs generated through our pipeline.
+As introduced in the previous section, GPT can also be used to generate virtual data in the corresponding format to enhance data quality.
 5 Experiment
 5.1 Quantitative Analysis
-Traditional NLP metrics such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015) evaluate generated text by comparing it to reference ground truth descriptions.
+Traditional NLP metrics such as BLEUPapineni et al.
+(2002), ROUGELin (2004) evaluate generated text by comparing it to reference ground truth descriptions.
 While effective for assessing surface-level language quality, these metrics fail to capture the richness of knowledge beyond the visual domain, such as audio-related information, character details, or contextual nuances.
@@
 These metrics evaluate fluency, knowledge integration (including both visual and audio elements), and narrative alignment, providing a more accurate assessment of the generated stories and better reflecting the advantages of our multi-modal framework.
-Language Fluency: To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story.
+Language Fluency: Fluency is used to evaluate the coherence and sentence repetition rate of generated explanatory texts.
+To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story.
 Repetition of triplets within the story is considered a sign of unnatural phrasing and redundancy, which negatively impacts fluency.
 A higher fluency score is assigned when the text flows smoothly, with minimal repetition and a coherent structure.
+We have made modifications to the formulas in the original paperYang et al.
+(2024).Fluency and repetition are opposite.
+In the chart, we use repetition to represent the fluency of language, with lower repetition indicating smoother language.
+Then repetition can be represented by Intra-story Repetition (ISR): $\text{ISR}(S)=\frac{1}{n(n-1)}\sum_{i\neq j}\frac{\left|W(s_{i})\cap W(s_{j})\right|}{\left|W(s_{i})\right|}$ where $n$ is the number of explanatory texts, and $W(s_{i})$ represents the vocabulary set of explanatory text $s_{i}$.
+The formula calculates the repetition rate of vocabulary within the explanatory texts, which should be minimized to ensure diversity and coherence in the text.
 Key-knowledge Relevance: For evaluating the relevance of key knowledge in the generated story, we define a knowledge base that includes both visual and audio information.
@@
 This expanded knowledge base allows us to assess the relevance and diversity of knowledge within the story.
-Specifically, we use two evaluation criteria: - Information Similarity (InfoSim): Measures the alignment between the knowledge points in the story and the knowledge repository.
+Specifically, we use two evaluation criteria: - Information Similarity (InfoSim)Yang et al.
+(2024): Measures the alignment between the knowledge points in the story and the knowledge repository.
 A higher similarity score indicates that the generated story effectively incorporates relevant knowledge.
-- Information Diversity (InfoDiverse): Evaluates the breadth of the knowledge used in the story.
+- Information Diversity (InfoDiverse)Yang et al.
+(2024): Evaluates the breadth of the knowledge used in the story.
 A higher diversity score indicates that the story incorporates a wide range of knowledge points, avoiding over-reliance on a small subset of information.
+The formula here is quite different from the original text.
+We have changed the normalization for time to the normalization for the number of keywords.
+That is to say, InfoDiverse calculates the proportion of keywords in the generated data that match the keywords in the reference data, ranging from 0 to 1.
+InfoSim $=\frac{1}{2|s_{i}|}\sum_{s_{i}}\left(\max_{k\in K}f_{k}^{T}f_{s_{i}}\right.$
+$\quad\left.+\frac{1}{|W(s_{i})|}\sum_{w\in W(s_{i})}\max_{k\in K}f_{k}^{T}f_{w}\right)$ (1)
+InfoDiverse $=\frac{1}{|K|}\left|\bigcup_{s_{i}}\{k_{t}\in K\mid\max_{w\in W(s_{i})\cup S_{i}}\right.$
+$\quad\left.f_{k_{t}}^{T}f_{w}>0.9\}\right|$ (2)
+where $W(s_{i})$ represents all words in sentence $s_{i}$.
+$f_{s_{i}}$, $f_{k}$, and $f_{w}$ refer to the normalized embeddings of sentence $s_{i}$, knowledge point $k$, and segmented word $w$, respectively.
 A well-rounded story should demonstrate both high information similarity and diversity, ensuring that relevant visual and audio knowledge are integrated in a meaningful way.
@@
 6 Conclusion
-Acknowledgments
-Appendix A Example Appendix
-This is an appendix.
-Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
-Internvideo2 61.98 38.00 11.24 0.257 0.295 0.108 0.0265 0.0154
-VideoLlava 15.94 38.18 13.64 0.376 0.275 0.113 0.0334 0.00474
-M2s(GPT4o) 10.24 30.10 59.11 0.175 0.148 0.0556 0.0128 1.25e-6
-M2s(llama) - - -
-M2s(Flan-T5) - - -
-Table 1: Performance of various models on Quantitative metrics
-Model Environment Language Emotion Character SRCI SACOR_T Overall GPT Human
-Internvideo2 1.0 1.0 0.0 0.6 0.2 0.600 0.56 -
-VideoLlava 1.1 1.2 0.0 0.9 0.9 0.559 0.8 -
-M2s(GPT4o) 2.7 2.3 1.0 1.7 1.7 0.735 1.9 -
-M2s(llama) - - - - - -
-M2s(Flan-T5) - - - - - -
-Table 2: Performance of various models on Qualitative metrics
-Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
-M2s(without emo&speed) 7.31 29.87 52.26 0.196 0.166 0.0651 0.0137 1.24e-6
-M2s(without ASR&emo&speed) 14.58 29.43 37.12 0.199 0.140 0.0626 0.0217 0.00365
-M2s(GPT4o) 10.24 28.63 83.30 0.175 0.148 0.0556 0.0128 1.25e-6
-Table 3: Performance of Ablation Study
+In this paper, we introduced MSBench, a comprehensive benchmark designed to evaluate the multi-modal story generation capabilities of MLLMs.
+We also proposed a robust MLLM pipeline, M2S, which outperforms leading models on the MSBench benchmark.
+Our extensive analyses provide valuable insights into the design of MLLMs for multi-modal story generation, particularly in scenarios enriched with additional information.
+Despite these advancements, our current approach has some limitations.
+We aim to address these in future work to enhance the evaluation framework and further refine the performance of MLLMs in complex, information-rich environments.
+Appendix A Appendix Tables
+Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
+Internvideo2 15.1
+VideoLlava 42.7
+Videochat2 8.1 38.0 55.0
+Internvideo2 + GPT4o 4.0 26.5 18.6 59.4 39.6 10.2 18.6 2.7
+VideoLlava + GPT4o 5.6 25.0 20.7 61.1 40.9 10.5 19.0 3.2
+Videochat2 + GPT4o 4.5 27.2 18.3 61.2 40.1 10.1 18.8 2.9
+Internvideo2 + A-GPT4o 5.6 25.5 17.6 67.0
+VideoLlava + A-GPT4o 6.0 25.2 21.9 65.3
+Videochat2 + A-GPT4o 12.4 27.3 53.6 52.8
+Table 1: Performance of various models with GPT4o as LLM on Quantitative metrics
+Model Environment Character Emotion Language SRCI Overall GPT Human
+Internvideo2 2.494 1.525 0.608 1.825 1.421 1.575
+VideoLlava 2.170 1.155 0.213 1.539 1.118 1.239
+Videochat2 3.086 1.878 0.921 2.243 1.809 1.987
+Internvideo2 + GPT4o 3.766 2.675 1.792 3.555 2.717 2.901
+VideoLlava + GPT4o 3.592 2.586 1.667 3.438 2.594 2.775
+Videochat2 + GPT4o 3.873 2.651 1.738 3.550 2.727 2.908
+Internvideo2 + A-GPT4o 3.581 2.574 1.659 3.433 2.585 2.767
+VideoLlava + A-GPT4o 3.427 2.491 1.607 3.195 2.485 2.641
+Videochat2 + A-GPT4o 3.492 2.568 1.673 3.090 2.625 2.689
+Table 2: Performance of various models with GPT4o as LLM on Qualitative metrics
+Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
+GPT4o 12.4 27.3 53.6 52.8 - - - -
+GPT3.5 13.5 26.3 25.5 54.0 38.2 8.6 17.6 1.9
+Doubao 5.4 27.7 16.0 56.8 32.6 7.0 15.7 0.8
+QWen 3.4 33.2 46.3 60.1 46.2 17.8 22.3 5.8
+QWen-7B-instruct 9.8 35.3 50.8 58.3 21.4 34.8 16.5 27.3
+mistral-instruct-7B-v1.0 8.1 8.3 7.5 3.5 4.7 1.8 6.3
+llama3-3B 22.1 - 7.9 16.6 8.6 21. 24.6 8.2
+lora_QWen-7B-instruct 7.6 33.1 50.8 58.2 52.4 21.7 25.1 9.0
+lora_mistral-instruct-7B-v1.0 7.3 33.0 50.1 58.0 52.2 21.7 25.0 8.9
+lora_llama3-3B 8.5 33.0 52.4 58.8 52.2 21.2 24.6 8.7
+Table 3: Performance of various LLMs on Quantitative metrics
+Model Environment Character Emotion Language SRCI Overall GPT Human
+GPT4o 3.492 2.558 1.673 3.090 2.625 2.689 -
+GPT3.5 2.831 1.809 1.077 2.299 1.815 1.966 -
+Doubao 2.998 1.906 1.025 2.233 1.878 2.008 -
+QWen 3.491 2.356 1.567 2.994 2.376 2.557 -
+QWen-7B-instruct 2.826 1.884 1.125 2.224 1.805 1.973 -
+mistral-instruct-7B-v1.0 0.360 0.229 0.124 0.297 0.213 0.245 -
+llama3-3B 0.294 0.187 0.084 0.261 0.134 0.192 -
+lora_QWen-7B-instruct 3.606 2.468 1.611 3.170 2.504 2.672 -
+lora_mistral-instruct-7B-v1.0 3.623 2.494 1.625 3.159 2.527 2.686 -
+lora_llama3-3B 3.560 2.431 1.577 3.106 2.457 2.626 -
+Table 4: Performance of various LLMs on Qualitative metrics